A 4.0520g sample of HCl, sp. Gr 11.18,

required 44.15ml of 0.9035Mof Sodium

Hydroxide in titration, compute for

percent purity.


1
Expert's answer
2021-05-11T05:51:12-0400

The balanced chemical equation for above is:


"NaOH+HCl\Rightarrow NaCl+H_2O"


Theoritical mass of "HCl" given = 4.0520 gm


Percent purity "= \dfrac{Experimental mass}{Theoritical mass} \times100"


Now we have to find experimental mass,which is calculated by the reaction.

For "NaOH"

M=0.9035 M

V=44.15 ml

so by the formula of molarity


"M=\dfrac {moles(n)}{volume(ml)} \times1000"


"0.9035= \dfrac{n_{NaOH}}{44.15}\times1000"


"n_{NaOH}= 0.04" moles


Now we will calculate the experimental moles of HCl

From the reaction:

1 mole of NaOH gave 1 mole of HCl

hence 0.04 moles of NaOH will gave 0.04 moles of HCl


Hence the experimental mass of HCl "=moles \times molar mass"

Experimental mass"=0.04\times36.5=1.46" gm


Now percent purity"=\dfrac{Experimental mass }{Theoriticalmass}\times100"


"=\dfrac{1.46}{4.0520}\times100=36.03" %